express as a mixed radical
a) square root 98
b) -5 square root 32
how do i solve these types of questions
thanks

Answer by stanbon(48535)

express as a mixed radical
a) square root 98 = sqrt(49*2) = 7sqrt(2)
----------------------------------------------
b) -5 square root 32 = -5sqrt(16*2) = -5*4sqrt(2) = -20sqrt(2)
======================
Cheers,
Stan H.